let X be a normed space and T belongs to B ( X).show that if a is eigen value of T then |a|<||T||
1
Expert's answer
2021-10-18T13:52:29-0400

From definition we have ||T(x)||"\le||T||\cdot ||x||" for any x"\in X"

Let a be eigenvalue of T, then "\exist x\in X,||x||\ne0" such that "T(x)=a\cdot x"

Therefore ||T(x)||=||a"\cdot" x||=|a|"\cdot" ||x||"\le ||T||\cdot| |x||"

Dividing both parts on value ||x||>0 we will have:

|a|"\le" ||T||, proof is done.

Statement |a|<||T|| is not true in the general case because if "X_0" ={"\lambda\cdot x,\lambda \in R" }, "T_{X0}(y)=a\cdot y,y\in X_0" we have ||T||=|a|.